can you give me the answer to this:

Identify the equation of direct variation given that y varies directly with x.
x is 84 when y is 6.


Answer by stanbon(75887)

Identify the equation of direct variation given that
y varies directly with x.
---
y = kx
---
Solve for "k" using "x is 84 when y is 6".
6 = k*84
k = 1/14
-----
Equation:
y = (1/14)x
====================
Cheers,
Stan H.
====================